OpenStudy (anonymous):

Find an approximate solution to the equation 10^(2x-1)=5. a. 1.0674 b. 0.8945 c. 2.7081 d. 6.6974

OpenStudy (anonymous):

start with \[2x-1=\log(5)\] then solve as you would any linear equation add 1, divide by 2 \[2x-1=\log(5)\] \[2x=\log(5)+1\] \[x=\frac{\log(5)+1}{2}\] now to get a decimal you will need a calculator

OpenStudy (anonymous):

I got 0.849485002168.

OpenStudy (anonymous):

yeah me too